# Tensor Completion via Integer Optimization

### v1

@@ 3.2 Relation between Gauge Norm and Tensor Rank
We next establish a relationship between tensor rank and our gauge norm, and we use this to argue that our norm is a meaningful constraint in the tensor completion problem. The underlying issue is related to the fact that the best low-rank approximation problem is not well-posed for tensors (de Silva & Lim, 2008), which is in sharp contrast to the case of nonnegative tensors for which the best low-rank approximation problem is well-posed (Qi et al., 2016).
@@
Assumption 3.5 (Regularity Condition).
@@
Proposition 3.7.
For any $\psi\in\Gamma$ that satisfies Assumption 3.5, we have $\|\psi\|_{\max}\leq\|\psi\|_{\pm}\leq\operatorname{rank}(\psi)\cdot\|\psi\|_{\max}$.
Proof.
The proof is similar to Proposition 2.4 of (Bugg et al., 2022). However, the right-side inequality requires the regularity assumption. Using the cp decomposition and the triangle inequality for norms,
$\begin{split}\|\psi\|_{\pm}\leq\textstyle\sum_{k=1}^{\operatorname{rank}(\psi)}\|\psi^{k}\|_{\pm}=\textstyle\sum_{k=1}^{\operatorname{rank}(\psi)}\|\psi^{k}\|_{\max},\\ \text{ where $\psi^{k}\in\mathcal{B}_{\infty}$}\end{split}$
The last equality follows by noting that $\|\psi^{k}\|_{\pm}=\|\psi^{k}\|_{\max}$ when $\psi^{k}\in\mathcal{B}_{\infty}$. Using $\|\psi^{k}\|_{\max}\leq\|\psi\|_{\max}$ from Assumption 3.5 gives the desired right-side inequality. ∎
@@ 4 Complexity Analysis of Norm
We show that calculating the norm $\|\ \cdot\|_{\pm}$ is np-hard. Despite this, it is still useful for tensor completion because it is defined using a convex polytope $\mathcal{C}_{1}$ whose vertices admit a convenient representation, as described in Proposition 3.4, that we will use to develop a practical algorithm.
Proposition 4.1 (Computational complexity).
@@
Proof.
Note that $\|\varphi\|_{\circ}=\sup\{|\langle\varphi,\psi\rangle|\ |\ \|\psi\|_{\pm}\leq 1\}=\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{C}_{1}\}$ is the dual norm for $\|\cdot\|_{\pm}$. The approximation of $\|\cdot\|_{\circ}$ can be reduced to approximation of the norm $\|\cdot\|_{\pm}$ in polynomial time from theorems 3 and 10 of (Friedland & Lim, 2016). Our main idea is to give a polynomial-time reduction of an np-Hard problem to an approximation of $\|\cdot\|_{\circ}$. In particular, we prove that calculating the $\infty,1$ subordinate matrix norm is polynomial-time reducible to $\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{S}_{1}\}$. Without loss of generality, assume $p=2$ and $d:=r_{1}=r_{2}$.
@@
From Proposition 1 of (Rohn, 2000), we have $\|W\|_{\infty,1}=\{\max\sum_{i,j}W_{ij}x_{i}y_{j}\>|\>x_{i},y_{j}\in\{-1,1\}\}$ for a matrix $W$. The decision version of the $\infty,1$ subordinate matrix norm for the special case of $M$-matrices can be written as:
@@
Clearly, setting $L=L^{\prime}$ and $\varphi=W$ is a valid polynomial time reduction. Since $\mathcal{C}_{1}=\operatorname{conv}(\mathcal{S}_{1})$ and $\langle\varphi,\psi\rangle$ is linear, $\|\varphi\|_{\circ}=\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{S}_{1}\}$. The result now follows since approximately solving any $\infty,p$ subordinate matrix norm, where $p\in[1,\infty)$, to arbitrary accuracy is np-hard (Hendrickx & Olshevsky, 2010). In particular, Theorem 5 of (Rohn, 2000) shows that it is an np-hard problem for $M$-matrices. ∎
The same proof of Corollary 3.2 in (Bugg et al., 2022) combined with the above result establishes np-completeness:
Corollary 4.2 (Bugg et al., 2022).
@@
Rademacher complexity, from computational learning theory, is used to characterize the richness of a class of functions (Bartlett & Mendelson, 2002; Srebro et al., 2010). Roughly speaking, function classes with lower Rademacher complexity can be learned using less samples. From the following proposition, one can check that the norm $\|\ \cdot\|_{\pm}$ has an exponentially smaller Rademacher complexity than the max and Frobenius norms for tensors.
Proposition 4.3 (Stochastic Complexity).
@@
Proof.
Rademacher complexity is computed using symmetric random variables $\sigma\in\{-1,+1\}$. It is for this reason that the class of functions $\mathcal{H}$, $-\mathcal{H}$, and $\mathcal{H}\cup-\mathcal{H}$ have the same Rademacher complexity. With this observation, the proof follows that of Proposition 3.3 of (Bugg et al., 2022). ∎
@@ 5.1 Complexity Analysis of Tensor Completion
By interpreting the tensor completion problem (2) as a convex aggregation problem (Nemirovski, 2000; Tsybakov, 2003; Lecué, 2013) for a finite set of functions, one can arrive at the generalization bound for the solution. Interestingly, we had got the same generalization bounds for the special case of nonnegative tensors (Bugg et al., 2022). We believe this is because the proof for nonnegative tensors did not exploit the non-negativity, and hence could have led to non-tight generalization bounds. For completeness, we state these statistical guarantees in the following two results:
Proposition 5.1 (Lecué, 2013).
@@
Corollary 5.2 (Bugg et al., 2022).
Suppose $\psi\in\Gamma$ is a tensor (satisfying Assumption 3.5) with $\operatorname{rank}(\psi)=k$ and $\|\psi\|_{\max}\leq\mu$. Under an additive noise model, if $(x\langle i\rangle,y\langle i\rangle)$ are independent and identically distributed with $|y\langle i\rangle-\varphi_{x\langle i\rangle}|\leq e$ almost surely and $\mathbb{E}y\langle i\rangle=\psi_{x\langle i\rangle}$. Then given any $\delta>0$, with probability at least $1-4\delta$ we have
@@
Unsurprisingly, the tensor completion problem (2) is np-hard, because approximating the norm $\|\cdot\|_{\pm}$ is np-hard and we had shown a polynomial-time reduction of the problem (2) to the np-hard weak membership problem in our previous work (Bugg et al., 2022).
Proposition 5.4.
@@
Proof.
The proof for this proposition is similar to the proof of Proposition 4.4 of (Bugg et al., 2022). ∎
@@ 5.2 Numerical Algorithm
The problem (2) is a convex optimization formulation despite being np-hard. This property enables the application of various first-order convex optimization algorithms. In particular, we use a variant of the Frank-Wolfe algorithm called Blended Conditional Gradients (BCG) (Braun et al., 2019). The main alternative approaches for iterative optimization seem to run into issues inherent to the implicit description of the feasible region; for instance, we do not have an effective barrier function available for interior point methods, nor do we have an efficient projection oracle that can be leveraged in projected gradient descent.
We find the BCG algorithm to be practically efficient for our problem since it requires a weak linear separation oracle leading to an early termination. When we design our weak separation oracle calls to an integer optimization problem, we explicitly define a tolerance for early termination. This works out well as integer optimization solvers usually discover near-optimal solutions fast and subsequently work hard to certify them. Further, the linear convergence of BCG is guaranteed when the feasible set is a polytope and the objective function is strictly convex. We construct a strictly convex problem by projecting the feasible space onto the set of unique, observed tensor entries like we did in (Bugg et al., 2022). Since the conditions for linear convergence are satisfied by our problem, we get an algorithm that terminates in a linear number of oracle steps.
@@
The BCG algorithm makes subsequent iterate updates based on the gradient of the objective function at the current iterate. It needs a weak separation oracle to find a new vertex that reduces a gradient-based linear objective. The weak separation oracle has the same requirements as in Algorithm 1 in (Bugg et al., 2022). The output of the oracle should either give a vertex that accomplishes separation, or a certificate that separation is not possible. We design our weak separation oracle using two algorithms, an integer optimization problem in (9) and an alternating maximization heuristic.
Since integer optimization is likely to be more computationally expensive than the alternating heuristic, we first try the latter a few times with different initializations. We extend our heuristic in Algorithm 2 of (Bugg et al., 2022) to general tensors, and explore solutions by toggling between $\{-1,1\}$ instead of $\{0,1\}$. It exploits the fact that the objective is multi-linear and minimizes the objective in different dimensions, going one by one. It runs in polynomial time and has been seen to speed up the computation in our simulations by reducing calls to the integer optimization solver. However, note that the heuristic is merely that and cannot give a certificate for the non-existence of separation.

### v2

@@ 3.2 Relation between Gauge Norm and Tensor Rank
We next establish a relationship between tensor rank and our gauge norm, and we use this to argue that our norm is a meaningful constraint in the tensor completion problem. The underlying issue is related to the fact that the best low-rank approximation problem is not well-posed for tensors desilva2008 , which is in sharp contrast to the case of nonnegative tensors for which the best low-rank approximation problem is well-posed qi2016 .
@@
Assumption 1 (Regularity Condition)
@@
Proposition 4
For any $\psi\in\Gamma$ that satisfies Assumption 1, we have $\|\psi\|_{\max}\leq\|\psi\|_{\pm}\leq rank(\psi)\cdot\|\psi\|_{\max}$.
Proof
The right-side inequality requires the regularity assumption. Using the cp decomposition and the triangle inequality for norms,
$\begin{split}\|\psi\|_{\pm}\leq\textstyle\sum_{k=1}^{rank(\psi)}\|\psi^{k}\|_{\pm}=\textstyle\sum_{k=1}^{rank(\psi)}\|\psi^{k}\|_{\max},\\ \text{ where $\psi^{k}\in\mathcal{B}_{\infty}$}\end{split}$
The last equality follows from $\|\psi^{k}\|_{\pm}=\|\psi^{k}\|_{\max}$ when $\psi^{k}\in\mathcal{B}_{\infty}$. Using $\|\psi^{k}\|_{\max}\leq\|\psi\|_{\max}$ from Assumption 1 gives the right-side inequality.
The proof of the left-side inequality is similar to Proposition 2.4 of bugg2022nonnegative . For any $\lambda\geq 0$, if $\psi\in\mathcal{S}_{\lambda}$, then by definition $\|\psi\|_{\max}=\lambda$. By the convexity of norms we have that: if $\psi\in\mathcal{C}_{\lambda}$, then $\|\psi\|_{\max}\leq\lambda$. This means that $\forall\>\lambda\geq 0$ we have $\mathcal{C}_{\lambda}\subseteq\mathcal{U}_{\lambda}:=\{\psi:\|\psi\|_{\max}\leq\lambda\}$, and thus $\inf\{\lambda\ |\ \psi\in\lambda\mathcal{U}_{1}\}\leq\inf\{\lambda\ |\ \psi\in\lambda\mathcal{C}_{1}\}$. But this is equivalent to $\|\psi\|_{\max}\leq\|\psi\|_{\pm}$, which proves the left-side inequality.
@@ 4 Complexity Analysis of Norm
We show that calculating the norm $\|\ \cdot\|_{\pm}$ is np-hard. Despite this, it is still useful for tensor completion because it is defined using a convex polytope $\mathcal{C}_{1}$ whose vertices admit a convenient representation, as described in Proposition 3, that we will use to develop a practical algorithm.
Proposition 5 (Computational complexity)
@@
Proof
Note that $\|\varphi\|_{\circ}=\sup\{|\langle\varphi,\psi\rangle|\ |\ \|\psi\|_{\pm}\leq 1\}=\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{C}_{1}\}$ is the dual norm for $\|\cdot\|_{\pm}$. The approximation of $\|\cdot\|_{\circ}$ can be reduced to approximation of the norm $\|\cdot\|_{\pm}$ in polynomial time from theorems 3 and 10 of friedland2016computational . Our main idea is to give a polynomial-time reduction of an np-Hard problem to an approximation of $\|\cdot\|_{\circ}$. In particular, we prove that calculating the $\infty,1$ subordinate matrix norm is polynomial-time reducible to $\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{S}_{1}\}$. Without loss of generality, assume $p=2$ and $d:=r_{1}=r_{2}$.
@@
From Proposition 1 of Rohn_2000 , we have $\|W\|_{\infty,1}=\{\max\sum_{i,j}W_{ij}x_{i}y_{j}\>|\>x_{i},y_{j}\in\{-1,1\}\}$ for a matrix $W$. The decision version of the $\infty,1$ subordinate matrix norm for the special case of $M$-matrices can be written as:
@@
Clearly, setting $L=L^{\prime}$ and $\varphi=W$ is a valid polynomial time reduction. Since $\mathcal{C}_{1}=conv(\mathcal{S}_{1})$ and $\langle\varphi,\psi\rangle$ is linear, $\|\varphi\|_{\circ}=\sup\{\langle\varphi,\psi\rangle\ |\ \psi\in\mathcal{S}_{1}\}$. The result now follows since approximately solving any $\infty,p$ subordinate matrix norm, where $p\in[1,\infty)$, to arbitrary accuracy is np-hard Hendrickx_Olshevsky_2010 . In particular, Theorem 5 of Rohn_2000 shows that it is an np-hard problem for $M$-matrices.
The same proof of Corollary 3.2 in bugg2022nonnegative combined with the above result establishes np-completeness:
Corollary 1 (bugg2022nonnegative )
@@
Rademacher complexity, from computational learning theory, is used to characterize the richness of a class of functions bartlett2002 ; srebro2010 . Roughly speaking, function classes with lower Rademacher complexity can be learned using less samples. From the following proposition, one can check that the norm $\|\ \cdot\|_{\pm}$ has an exponentially smaller Rademacher complexity than the max and Frobenius norms for tensors.
Proposition 6 (Stochastic Complexity)
@@
Proof
The proof is similar to Proposition 3.3 of bugg2022nonnegative . By definition, $\mathsf{R}(\mathcal{C}_{\lambda})=\mathbb{E}_{\sigma}(\sup_{\Psi\in\mathcal{C}_{\lambda}}\textstyle\frac{1}{n}|\sum_{i=1}^{n}\sigma_{i}\cdot\psi_{x\langle i\rangle}|)$ and $\mathsf{W}(\mathcal{C}_{\lambda})=\sup_{X}\mathsf{R}(\mathcal{C}_{\lambda})$, where $\sigma_{i}$ are independent and identically distributed (i.i.d.) Rademacher random variables (i.e., $\sigma_{i}=\pm 1$ with probability $\frac{1}{2}$) bartlett2002 ; srebro2010 . Hence, $\mathsf{R}(\mathcal{C}_{\lambda})\leq\mathsf{W}(\mathcal{C}_{\lambda})$. Recall that $\mathcal{C}_{\lambda}=conv(\mathcal{S}_{\lambda})$ by Proposition 1. This means $\mathsf{W}(\mathcal{C}_{\lambda})=\mathsf{W}(\mathcal{S}_{\lambda})$ ledoux1991 ; bartlett2002 . Next, observe that
$\mathsf{W}(\mathcal{C}_{\lambda})$ $=\mathsf{W}(\mathcal{S}_{\lambda})\textstyle=\sup_{X}\mathbb{E}_{\sigma}\big{(}\sup_{\psi\in\mathcal{S}_{\lambda}}\textstyle\frac{1}{n}\big{|}\sum_{i=1}^{n}\sigma_{i}\cdot\psi_{x\langle i\rangle}\big{|}\big{)}$ (7)
$\textstyle=\sup_{X}\mathbb{E}_{\sigma}\big{(}\max_{\psi\in\mathcal{S}_{\lambda}}\textstyle\frac{1}{n}\cdot\sum_{i=1}^{n}\sigma_{i}\cdot\psi_{x\langle i\rangle}\big{)}\leq\sup_{X}r\sqrt{2\log\#\mathcal{S}_{\lambda}}/n$
where in the second line since the set $\mathcal{S}_{\lambda}$ is finite, we used the Finite Class Lemma massart2000 with $r=\max_{\psi\in\mathcal{S}_{\lambda}}\textstyle\sqrt{\sum_{i=1}^{n}(\psi_{x\langle i\rangle})^{2}}\leq\lambda\sqrt{n}$, and replaced the supremum with a maximum. This inequality on $r$ is because $\mathcal{S}_{\lambda}$ consists of tensors whose entries are from $\{-\lambda,\lambda\}$. Thus $\mathsf{W}(\mathcal{C}_{\lambda})\leq\lambda\sqrt{(2\log 2)\cdot\rho/n}\leq\lambda\cdot 2\sqrt{\rho/n}$.
@@ 5.1 Complexity Analysis of Tensor Completion
By interpreting the tensor completion problem (2) as a convex aggregation problem nemirovski2000topics ; tsybakov2003optimal ; lecue2013empirical for a finite set of functions, one can arrive at the generalization bound for the solution. Interestingly, these generalization bounds are the same in the special case of nonnegative tensors bugg2022nonnegative . We believe this is because the proof for nonnegative tensors did not exploit the non-negativity, and hence could have led to non-tight generalization bounds. For completeness, we state these statistical guarantees in the following two results:
Proposition 7 (lecue2013empirical )
@@
Corollary 2 (bugg2022nonnegative )
Suppose $\psi\in\Gamma$ is a tensor (satisfying Assumption 1) with $rank(\psi)=k$ and $\|\psi\|_{\max}\leq\mu$. Under an additive noise model, if $(x\langle i\rangle,y\langle i\rangle)$ are independent and identically distributed with $|y\langle i\rangle-\varphi_{x\langle i\rangle}|\leq e$ almost surely and $\mathbb{E}y\langle i\rangle=\psi_{x\langle i\rangle}$. Then given any $\delta>0$, with probability at least $1-4\delta$ we have
@@
Unsurprisingly, the tensor completion problem (2) is np-hard, because approximating the norm $\|\cdot\|_{\pm}$ is np-hard and there is a polynomial-time reduction of the problem (2) to the np-hard weak membership problem bugg2022nonnegative .
Proposition 8
@@
Proof
The proof for this proposition is similar to the proof of Proposition 4.4 of bugg2022nonnegative . Define the ball of radius $\delta>0$ centered at a tensor $\psi$ to be $B(\psi,\delta)=\{\varphi:\|\varphi-\psi\|_{F}\leq\delta\}$. Next define $W(\mathcal{C}_{1},\delta)=\bigcup_{\psi\in\mathcal{C}_{1}}B(\psi,\delta)$ and $W(\mathcal{C}_{1},-\delta)=\{\psi\in\mathcal{C}_{1}:B(\psi,\delta)\subseteq\mathcal{C}_{1}\}$. The weak membership problem for $\mathcal{C}_{1}$ is that given a tensor $\psi$ and a $\delta>0$ decide whether $\psi\in W(\mathcal{C}_{1},\delta)$ or $\psi\notin W(\mathcal{C}_{1},-\delta)$. Theorem 10 of friedland2016computational shows that approximation of the norm $\|\cdot\|_{\pm}$ is polynomial-time reducible to the weak membership problem for $\mathcal{C}_{1}$. Since Proposition 5 shows that approximation of the norm $\|\cdot\|_{\pm}$ is NP-hard, the result follows if we can reduce the weak membership problem to (2).
Suppose we are given inputs $\psi$ and $\delta$ for the weak membership problem. Choose $x\langle i\rangle$ for $i=1,\ldots,\pi$ such that each element in $\mathcal{R}$ is enumerated exactly once. Next choose $y\langle i\rangle=\psi_{x\langle i\rangle}$ and $\lambda=1$. Finally, note if we solve (2) and the minimum objective value is less than or equal to $\delta$, then we have $\psi\in W(\mathcal{C}_{1},\delta)$; otherwise, we have $\psi\notin W(\mathcal{C}_{1},-\delta)$. The result now follows since this was the desired reduction.
@@ 5.2 Numerical Algorithm
The problem (2) is a convex optimization formulation despite being np-hard. This property enables the application of various first-order convex optimization algorithms. In particular, we use a variant of the Frank-Wolfe algorithm called Blended Conditional Gradients (BCG) braun2019blended . The main alternative approaches for iterative optimization seem to run into issues inherent to the implicit description of the feasible region; for instance, we do not have an effective barrier function available for interior point methods, nor do we have an efficient projection oracle that can be leveraged in projected gradient descent.
@@
The BCG algorithm makes subsequent iterate updates based on the gradient of the objective function at the current iterate. It needs a weak separation oracle to find a new vertex that reduces a gradient-based linear objective. The weak separation oracle is given in Algorithm 1 bugg2022nonnegative . The output of the oracle should either give a vertex that accomplishes separation, or a certificate that separation is not possible. We design our weak separation oracle using two algorithms, an integer optimization problem in (12) and an alternating maximization heuristic.
Since integer optimization is likely to be more computationally expensive than the alternating heuristic, we first try the latter several times with different randomized initializations. The heuristic adapts Algorithm 2, where solutions are explored by toggling between ±$\pm 1$ bugg2022nonnegative . The following (nonlinear) separation objective is minimized over binary $\theta$, given a BCG-generated parameter $c$:
$z_{M}(\theta):=\sum_{x\in\mathcal{R}}\langle c_{x},\psi_{x}-\textstyle\lambda\cdot\prod_{k=1}^{p}\theta_{x_{k}}\rangle.$ (11)
  Input: linear objective $c\in\mathbb{R}^{r_{1}\times\cdots\times r_{p}}$, point $\psi\in\mathcal{C}_{\lambda}$, accuracy $K\geq 1$, gap estimate $\Phi>0$, norm bound $\lambda$
1  Output: Either (1) vertex $\varphi\in\mathcal{S}_{\lambda}$ with $\langle c,\psi-\varphi\rangle\geq\Phi/K$, or (2) false:
$\langle c,\psi-\varphi\rangle\leq\Phi$ for all $\varphi\in\mathcal{C}_{\lambda}$
Algorithm 1 Weak Separation Oracle for $\mathcal{C}_{\lambda}$
1  Input: linear objective $c\in\mathbb{R}^{r_{1}\times\cdots\times r_{p}}$, point $\psi\in\mathcal{C}_{\lambda}$, norm bound $\lambda$, incumbent
solution $\hat{\theta}\in\mathcal{S}_{\lambda}$, objective function $z_{M}(\theta):=\sum_{x\in\mathcal{R}}\langle c_{x},\psi_{x}-\textstyle\lambda\cdot\prod_{k=1}^{p}\theta_{x_{k}}\rangle$.
  Output: Best known solution $\theta$
  $\theta\leftarrow\hat{\theta}$
  $z\leftarrow z_{M}(\theta)$
  for $i=1$ to $p$ do
     for $k=1$ to $r_{i}$ do
        $\theta^{(i)}_{k}\leftarrow-1\cdot\theta^{(i)}_{k}$
        if $z_{M}(\theta)>z$ then
           $z\leftarrow z_{M}(\theta)$
        else
           $\theta^{(i)}_{k}\leftarrow-1\cdot\theta^{(i)}_{k}$
        end if
     end for
  end for
Algorithm 2 Alternating Maximization
The heuristic considers one dimension at a time, setting the corresponding entries of $\theta$ to optimize within the local neighbourhood defined by the given dimension (which can be done by simply considering the signs of $c$). It runs in polynomial time (linear in the size of $\theta$) and has been seen to speed up the computation in our simulations by reducing calls to the integer optimization solver. However, note that the heuristic is merely that and cannot in general give a certificate for the non-existence of separation, which requires global optimization. Indeed, even in the case of a matrix ($p=2$), $z_{M}$ represents a generic form of NP-hard Quadratic Unconstrained Binary Optimization (QUBO) (see, e.g. lewis2017quadratic ).
@@
Note that, in principle, for some binary variable $q\in\{-1,1\}$, one can directly branch on the disjunction:
$q=-1\lor q=1.$
However, in implementation we choose instead to reformulate via 0-1 auxiliary variables, $\sigma:=(q+1)/2$, as enforcing $\sigma\in\{0,1\}$ obviates the need for imposing $q\in\{-1,1\}$. Integer optimization solvers typically can only handle 0-1 binary variables by default instead of $\{-1,1\}$ due to the more extensive set of techniques developed for 0-1 variables.
@@
Proposition 9
Using BCG approach, the convex optimization formulation of the tensor completion problem in (2) can be solved in a linear number of calls to the Weak Separation Oracle in Algorithm 1.
Proof
The linear convergence of BCG is guaranteed when the feasible set is a polytope and the objective function is strictly convex. Note that the feasible set $\|\psi\|_{\pm}\leq\lambda$ in Problem (2) is a polytope. Further, it is easy to construct a strictly convex problem by projecting the feasible space onto the set of unique, observed tensor entries bugg2022nonnegative . Specifically, we use the equivalent reformulation in which we change the feasible set from $\{\psi:\|\psi\|_{\pm}\leq\lambda\}=C_{\lambda}$ to $\mathrm{Proj}_{U}(\mathcal{C}_{\lambda})$ where the projection is done over the unique indices specified by the set $U$. Since the conditions for linear convergence are satisfied by our problem, our algorithm terminates in a linear number of oracle calls to Algorithm 1.
In fact, we find the BCG algorithm to be practically efficient for our problem since the weak linear separation oracle can accept (sufficiently good) suboptimal solutions that may be found at early termination of a global solver. When we design our weak separation oracle calls to an integer optimization problem, we explicitly define a tolerance for early termination. This works out well as integer optimization solvers usually discover near-optimal solutions fast and subsequently work hard to certify them.